# GUing: A Mobile GUI Search Engine using a Vision-Language Model

### v2

@@ 8.3.2. External Validity
The selection of queries is crucial for the evaluation, since the search performance may vary depending on the query. For example, a GUI search engine that exhibits satisfactory results for the health and fitness domain may not necessarily produce the same outcome for the finance or education domains. It is thus imperative to consider varying domains and apps in the evaluation. The study of Kolthoff et al. (Kolthoff et al., 2023) provides a gold standard of queries and corresponding GUIs. However, this dataset has already been exhausted for the training of RaWi, which makes it unsuitable for our comparative study. Given the lack of alternative query datasets, we searched for articles published by companies specialised in mobile app development. This decision could potentially introduce a selection bias. The app features extracted from the articles span ten distinct domains (Terekhov, 2023), plus innovative features from multiple additional domains (Hughart, 2023). Therefore, we think that these sets are representative enough for the purpose of our evaluation. Nevertheless, reproducing our study with additional queries, scenarios, and evaluators would certainly further strengthen the generalisablity of its results.
@@ 9. Conclusion and Future Work
In this paper, we presented GUing, a GUI search engine, built upon GUIClip, a novel vision-language foundation model for the domain of mobile apps. The GUIClip model was trained using SCapRepo, a comprehensive dataset comprised of 135k screenshot-caption pairs created by actual app experts. In addition, we presented ScreenRepo, a dataset that serves as a large repository for GUing. SCapRepo and ScreenRepo were created through an automated pipeline, enabling an easy update to the datasets, e.g. when new screen designs emerge.
Our evaluation, including a benchmarking on various datasets and a manual assessment of search results, indicates that GUing outperforms state-of-the-art approaches for GUI retrieval. Moreover, the evaluation highlights the critical role that the SCapRepo and ScreenRepo datasets play in achieving the top performance of our vision-language models: not only for GUI retrieval but also for GUI classification and sketch-to-GUI retrieval in the mobile app domain. From here, there are several follow-up future directions:

### v3

@@ 7.3.2. External Validity
We discuss the threats to external validity for the evaluation of GUing and GUIClip.
GUing: Query Selection for Search Engine Evaluation. The selection of queries is crucial for the evaluation, since the search performance may vary depending on the query. For example, a GUI search engine that exhibits satisfactory results for the health and fitness domain may not necessarily produce the same outcome for the finance or education domains. It is thus imperative to consider varying domains and apps in the evaluation. The study of Kolthoff et al. (Kolthoff et al., 2023) provides a gold standard of queries and corresponding GUIs. However, this dataset has already been exhausted for the training of RaWi, which makes it unsuitable for our comparative study. Given the lack of alternative query datasets, we searched for articles published by companies specialised in mobile app development. This decision could potentially introduce a selection bias. The app features extracted from the articles span ten distinct domains (Terekhov, 2023), plus innovative features from multiple additional domains (Hughart, 2023). Therefore, we think that these sets are representative enough for the purpose of our evaluation. Nevertheless, reproducing our study with additional queries, scenarios, and evaluators would certainly further strengthen the generalisability of its results.
GUIClip: Bias of Datasets for Model Evaluation. Enrico contains 1460 screenshots for 20 categories, while Swire contains 3551 sketch-screenshot pairs. Compared with the large scale of SCapRepo, these datasets used to evaluate GUIClip for other GUI-related tasks are relatively small. This may introduce potential bias into evaluation. To address this, we manually reviewed these two datasets and found that they encompass diverse domains. This suggests that GUIClip is effective for handling various GUIs across multiple domains. However, to strengthen the external validity of the findings, further evaluation on additional datasets is necessary, ideally with a broader range of domains, apps, and designs.
@@ 8. Conclusion and Future Work
In this paper, we presented GUing, a GUI search engine, built upon GUIClip, a novel vision-language foundation model for the domain of mobile apps. The GUIClip model was trained using SCapRepo, a comprehensive dataset comprised of 135k screenshot-caption pairs created by actual app experts. In addition, we presented ScreenRepo, a dataset that serves as a large repository for GUing. SCapRepo and ScreenRepo were created through an automated pipeline, enabling an easy update, e.g. when new screen designs emerge.
Our evaluation, including a benchmarking on various datasets and a manual assessment of search results, indicates that GUing outperforms state-of-the-art approaches for GUI retrieval. Moreover, the evaluation highlights the critical role that the SCapRepo and ScreenRepo datasets play in achieving the top performance of our vision-language model: not only for GUI retrieval but also for GUI classification and sketch-to-GUI retrieval in the mobile app domain. From here, there are several follow-up future directions: